OpenStudy (anonymous):

f(x)=x^2 .... Find f(1.1)-f(1)/(1.1-1)

OpenStudy (anonymous):

Pls help me!

OpenStudy (unklerhaukus):

\[f(x)=x^2\\ \\f(1)=(1)^2\\f(1.1)=(1.1)^2\] \[\frac{f(1.1)-f(1)}{1.1-1}=\frac{(1.1)^2-(1)^2}{1.1-1}=\]